find (a)2v + u, (b)u -3v, and (c)5u - v in the equation of 11. u = <-2, -3>, v = <-1,-10>
So far I have,
u=(-2,3)
v=(-1,-10)
2v+u=(-2,-20)+(-1,-10)
2v+u=(-2+(-1),-20+(-10))
2v+u=(-3,-30)
but need to label (a), (b), and (c). Ty.

Answer by themasterofcircuits(27)

u=<-2,3> and v=<-1,-10>
a.) 2v+u:
2*<-1,-10>+<-2,-3>=<-2-2,-20-3> = [<-4,-23>]
b.) u-3v
<-2,-3>-3*<-1,-10>=<-2+3,-3+30> = [<1,27>]
c.) 5u-v
5*<-2,-3>-<-1,-10>=<-10+1,-15+10>= [<-9,-5>]